Norma finished her math assignment in 1/3 hours. Then she completed her chemistry assignment in 1/8 hours. What was the total amount of time Norma spent doing these two assignments?
1
Expert's answer
2019-06-10T10:31:52-0400

Total amount of time: "t=\frac{1}{3}+\frac{1}{8}=\frac{8}{24}+\frac{3}{24}=\frac{11}{24}\;hrs\;=27.5\;min."